A $100,000 mortgage is to be amortized by making monthly payments for 20 years. Interest is 5.1% compounded semi-annually for a six-year term.

(a)

(b)

(c)

Compute the size of the monthly payment.

Determine the balance at the end of the six-year term.

If the mortgage is renewed for a six-year term at 5% compounded semi-annually, what is the size of the monthly payment for the renewal term?

(a) The size of the monthly payment is $

(Round the final answer to the nearest cent as needed. Round all intermediate values to six decimal places as needed.)

(b) The balance at the end of the six-year term is $

(Round the final answer to the nearest cent as needed. Round all intermediate values to six decimal places as needed.)

(c) The size of the monthly payment for the renewal term is $

(Round the final answer to the nearest cent as needed. Round all intermediate values to six decimal places as needed.)


1
Expert's answer
2021-12-09T12:07:18-0500

solutions according to the formula:

a)"A =\frac{ P (\frac{r}{2})\times(1 + \frac{r}{2})^n}{(1 + \frac{r}{2})^n- 1}=\frac{100000(\frac{0.051}{2})\times(1 + \frac{0.051}{2})^{40}}{(1 + \frac{0.051}{2})^{40}-1}=4017.24"


b)"4017.24\times40-4017.24\times6=136586.17"


c)"A =\frac{ P (\frac{r}{2})\times(1 + \frac{r}{2})^n}{(1 + \frac{r}{2})^n- 1}=\frac{100000(\frac{0.05}{2})\times(1 + \frac{0.05}{2})^{52}}{(1 + \frac{0.05}{2})^{52}-1}=3457.44"